OpenStudy (anonymous):

Solve the triangle. A = 50°, b = 13, c = 6 choices: a ≈ 10.2, C ≈ 26.3, B ≈ 103.7 a ≈ 14, C ≈ 26.3, B ≈ 103.7 No triangles possible a ≈ 14, C ≈ 30.3, B ≈ 99.7

OpenStudy (anonymous):

@lgbasallote

OpenStudy (amistre64):

law of sines and cosines

OpenStudy (lgbasallote):

ugh. what he/she said

OpenStudy (anonymous):

i dont get it tho!!! can you show me.

OpenStudy (amistre64):

how is the law of cosine defined?

OpenStudy (anonymous):

im not sure. im no good at math

OpenStudy (amistre64):

well, this is why they write down the formula in the pages of the textbooks ....

OpenStudy (amistre64):

it a generalization of the pythag thrm; do you recall the pythag thrm by chance?

OpenStudy (anonymous):

no

OpenStudy (amistre64):

then you are clearly not ready to tackle this problem

OpenStudy (anonymous):

thats why i am clearly here.

mathslover (mathslover):

do you know about trigonometry?

OpenStudy (anonymous):

nope. im taking pre-calc

OpenStudy (amistre64):

precalc IS trigonometry

OpenStudy (anonymous):

oh well then still no.

mathslover (mathslover):

Have you just started learning that?

OpenStudy (anonymous):

yes

mathslover (mathslover):

After all one 'yes'

mathslover (mathslover):

http://www.youtube.com/watch?v=F21S9Wpi0y8

OpenStudy (anonymous):

i just started, and im doing my first assignment. so i needed help.

OpenStudy (anonymous):

im in school we dont have access to youtube

mathslover (mathslover):

oh! then did you try to ask to your teacher?

OpenStudy (amistre64):

your first assignment is to play with the law of sines and cosines? they usually tend to build up to those things, or at least start you off teaching the trig functions

OpenStudy (anonymous):

im doing virtual school, but its taken in school. shes not available right now,.

OpenStudy (anonymous):

why are you questioning me so much? i just want help on this problem.

OpenStudy (amistre64):

this IS the help. http://openstudy.com/code-of-conduct

mathslover (mathslover):

We an not allowed to directly ANSWER problems. --- You must cooperate us by telling what You tried

OpenStudy (anonymous):

your not that funny. im not looking for direct answers. im looking for something called H-E-L-P. you obviously dont know how to administor that.

OpenStudy (amistre64):

good luck then

OpenStudy (anonymous):

goodluck in learning how to help people.

mathslover (mathslover):

spaz daw actually to start with this you must know about sines, cosines

mathslover (mathslover):

that you don't

OpenStudy (anonymous):

hence my resoning for coming to this site.

OpenStudy (lgbasallote):

virtual schools love skipping fundamentals @mathslover

OpenStudy (lgbasallote):

just show him an example/demonstration of the cosine law...im sure he'll pick it up

OpenStudy (anonymous):

*she

OpenStudy (lgbasallote):

my keyboard ha trouble preing that letter

OpenStudy (anonymous):

lol

OpenStudy (lgbasallote):

i would love to give you a demnstration of how cosine law works...but i cannot remember it anymore. I only know sine law. sorry

OpenStudy (amistre64):

the pythag thrm is a special case of the law of cosines: pythag thrm: c^2 = a^2 + b^2 the law of cosine generalizes this to any SAS triangle:\[c^2=a^2+b^2-2ab~cos(A)\] when A = 90 degree, the end term goes to zero and we end up with the pythag thrm

OpenStudy (amistre64):

for a proof ill use this drawing 

OpenStudy (amistre64):

by constructing this into a 90 degree triangle, we can compare this with the better known pythag thrm